The program continues to tell me my answers are wrong but I am not sure why. See the question and my answers below:liabilities and stockholders equity of The following table presents the long-term Information Control Corp. one year ago: $
Answer #1

Here, some rectification in your answer based on the concept of issue of new shares and debt .

  • The long-term debt account will increase by $35.6 million,
  • common stock account should be increase by $10.6 million as 10.6 million shares with $1 par value
  • additional capital surplus =59.2-10.6 = $48.6 million
  • Addition to retained earnings =Net income- dividend = $10.2 -2.6= $7.6 million


Long-term debt $ 101,200,000
Total long-term debt $ 101,200,000
Shareholders equity
Preferred stock $ 4,060,000
Common stock ($1 par value) $ 26,200,000
Capital surplus $ 94,200,000
Accumulated retained earnings $ 143,200,000
Total equity $ 267,660,000
Total Liabilities & Equity $ 368,860,000